OpenStudy (anonymous):

What is the simplified form of THE SQUARE ROOT of 48n^9 ?

OpenStudy (hoblos):

sqrt(48n^9) = sqrt(16*3*n*n^8) = 4n^4 sqrt(3n)

OpenStudy (anonymous):

THANKS!!

OpenStudy (hoblos):

my pleasure XD